### v1

# Panacea [Uncaptioned image]: Mitigating Harmful Fine-tuning for Large Language Models via Post-fine-tuning Perturbation

Abstract
Harmful fine-tuning attack introduces significant security risks to the fine-tuning services. Main-stream defenses aim to vaccinate the model such that the later harmful fine-tuning attack is less effective. However, our evaluation results show that such defenses are fragile– with a few fine-tuning steps, the model still can learn the harmful knowledge. To this end, we do further experiment and find that an embarrassingly simple solution– adding purely random perturbations to the fine-tuned model, can recover the model from harmful behaviors, though it leads to a degradation in the model’s fine-tuning performance. To address the degradation of fine-tuning performance, we further propose Panacea, which optimizes an adaptive perturbation that will be applied to the model after fine-tuning. Panacea maintains model’s safety alignment performance without compromising downstream fine-tuning performance. Comprehensive experiments are conducted on different harmful ratios, fine-tuning tasks and mainstream LLMs, where the average harmful scores are reduced by up-to 21.5%, while maintaining fine-tuning performance. As a by-product, we analyze the adaptive perturbation and show that different layers in various LLMs have distinct safety coefficients. Source code available at https://github.com/w-yibo/Panacea.
Machine Learning, ICML
1 Tsinghua University; 2 Shenzhen Campus of Sun Yat-sen University; 3 Didichuxing Co. Ltd; 4 Nanyang Technological University. Correspondence to: Li Shen <mathshenli@gmail.com>.
1 Introduction
Fine-tuning-as-a-service (OpenAI, ) is a popular business service to enhance model’s performance for customized datasets, domain-specific tasks, and private needs. However, recent studies (Qi et al., 2023; Zhan et al., 2023; Yang et al., 2023; Yi et al., 2024a; Li et al., 2024a; Ye et al., 2024) identify a safety issue, the harmful fine-tuning attack (Figure 1), where the model’s safety alignment is compromised when the fine-tuning dataset contains harmful data, even a small amount of harmful data can introduce significant security vulnerabilities. Moreover, harmful fine-tuning is often unintentional, as datasets may contain latent unsafe data that is difficult for users to detect. Since service providers are responsible for the harmful outputs generated by the model, there is a clear need for more effective solutions.
Figure 1: The harmful fine-tuning attack for fine-tuning-as-a-service scenarios. In the alignment stage, pretrained LLMs are aligned using alignment data to produce aligned LLMs. In the fine-tuning stage, aligned LLMs are further fine-tuned using fine-tuning data, which may contain harmful data, potentially leading to unsafe fine-tuned models.
Vaccine (Huang et al., 2024e) and RepNoise (Rosati et al., 2024a) are two representative defenses against the harmful fine-tuning attack. Vaccine improves the aligned model’s resistance towards harmful knowledge learning by solving a mini-max problem, while RepNoise aims to erase the harmful knowledge of the aligned model by perturbing the hidden representation given harmful data. However, our evaluation results show that with more fine-tuning steps, the vaccinated model produced by the two methods are still suffering from the negative effect of harmful fine-tuning attack – their harmful score are significantly increased, though with a slower rate compared to the baseline without defense.
Based on the above finding, it seems that the learning of harmful knowledge cannot be sufficiently suppressed before fine-tuning. From another angle, it may be worthwhile to consider a mitigation approach to the problem after fine-tuning. We start our exploration by a rather naive defense– adding purely random post-fine-tuning perturbation to the fine-tuned model. Our evaluation results surprisingly demonstrate that random perturbation can recover the model from harmful behavior, showing that such a naive method could be a potential defense. However, our subsequent evaluation shows that this method significantly degrades the fine-tuning performance of the model, indicating that such a method cannot strike a good balance between recovering from harmful behavior and maintaining fine-tuning performance. To this end, a subsequent question is that:
How to add post-fine-tuning perturbation to the fine-tuned model, such that it can be recovered from harmful behavior without hurting downstream task performance much?
Driven by this question, we propose Panacea (Figure 2), an iterative gradient-based method to iterately search for the post-fine-tuning perturbation. Panacea aims to solve a max-maximize optimization problem, such that the added perturbation can maximally increase the model’s harmful loss, ensuring that it can effectively recover the model from harmful behaviors. The experiment results show that for different harmful ratios during fine-tuning, our method’s average harmful score is reduced by 10.4% up to 21.5%, while fine-tuning performance improves by 0.3% over the standard alignment method. The ablation study on the adaptive perturbation show that it can reduce harmful scores by up to 23.2%, while maintaining competitive fine-tuning performance. The visualization experiments also reveal different layers in various LLMs have distinct safety coefficients, consistent with previous findings and providing additional evidence for layer-wise safety research.
The main contributions of this work contain three parts:
1. We find that adding purely random perturbations to the fine-tuned models could recover the model from harmful behavior, but does cause a loss of fine-tuning performance.
2. To mitigate harmful fine-tuning while maintaining the fine-tuning performance, we propose Panacea, a post-fine-tuning solution that formulates a max-max optimization problem, where the optimized perturbation maximally increases the harmful loss.
3. Experiments evaluate Panacea across diverse settings, demonstrating its effectiveness and generalization, while visualizations reveal safety coefficients across LLM layers.
2 Related Work
Safety Alignment. The safety alignment requires the model to output content that is both helpful and harmless, and to be able to output a refusal answer when given harmful instructions. Existing methods typically rely on supervised fine-tuning (SFT), RLHF (Ouyang et al., 2022), and variations of RLHF (e.g., PPO, DPO) (Bai et al., 2022; Wu et al., 2023; Dong et al., 2023; Rafailov et al., 2023; Yuan et al., 2023; Huang et al., 2021). These methods construct a safety-aligned dataset, and recent approaches focus on enhancing and better utilizing the aligned dataset (Liu et al., 2023a, b; Ye et al., 2023; Tekin et al., 2024).
Figure 2: Post-fine-tuning perturbation. The fine-tuned model exhibits a high harmful score (HS:$\downarrow$). Adding random perturbation reduces the harmful score but also decreases fine-tuning accuracy (FA:$\uparrow$). In contrast, incorporating our post-fine-tuning perturbation (see Algorithm 1) effectively lowers the harmful score while maintaining fine-tuning performance.
Harmful Fine-tuning. Fine-tuning-as-a-service becomes a mainstream method for LLMs API providers. Recent studies (Qi et al., 2023; Zhan et al., 2023; Yang et al., 2023; Li et al., 2024a; Ye et al., 2024; Lermen et al., 2023; He et al., 2024; Halawi et al., 2024; Chen et al., 2024; Hawkins et al., 2024; Poppi et al., 2024) show that LLMs trained with safety alignment can be jail-broken when fine-tuned on a dataset with a small amount of harmful data. In such cases, the model fails to refuse harmful instructions and outputs harmful responses. Many works (Rosati et al., 2024c; Leong et al., 2024; Anonymous, 2025c; Guo et al., 2024; Qi et al., 2024b; Rosati et al., 2024a) focus on analyzing the mechanisms of different harmful fine-tuning attacks. (Peng et al., 2024) proposes new safety metrics to evaluate harmful fine-tuning risk and (Rosati et al., 2024a) explores the safety risks when learning with reinforcement learning. Existing defenses can be categorized into three main categories (Huang et al., 2024c), i) Alignment stage solutions (Huang et al., 2024e; Rosati et al., 2024b; Tamirisa et al., 2024; Liu et al., 2024c; Pan et al., ; Huang et al., 2024b; Anonymous, 2025a; Liu et al., 2024a), ii) Fine-tuning stage solutions (Mukhoti et al., 2023; Bianchi et al., 2023; Zong et al., 2024; Wei et al., 2024; Huang et al., 2024d; Wang et al., 2024; Lyu et al., 2024; Eiras et al., 2024; Zhang et al., 2024; Li et al., 2024b; Shen et al., 2024; Anonymous, 2025b; Li et al., 2025; Du et al., 2024a; Choi et al., 2024; Luo et al., 2024; Qi et al., 2024a), iii) Post-fine-tuning stage solutions (Casper et al., 2024; Yi et al., 2024c; Du et al., 2024b; Hsu et al., 2024; Huang et al., 2024a; Zhu et al., 2024; Liu et al., 2024b; Wu et al., 2024; Yi et al., 2024b). The proposed method in this paper is applied in the post-fine-tuning stage, aiming to restore safety alignment without sacrificing fine-tuning performance.
Mainstream defenses focusing on the alignment stage lack sufficient durability against harmful fine-tuning (Qi et al., 2024b), motivating exploration of the post-fine-tuning stage. Existing post-fine-tuning solutions typically add perturbations based on prior knowledge, such as a safety subspace (Hsu et al., 2024) or safety-critical parameters (Zhu et al., 2024; Yi et al., 2024b). In contrast, Panacea optimizes an adaptive perturbation during fine-tuning without relying on prior knowledge.
3 Revisiting Alignment Defenses
3.1 Problem Setup
Harmful Fine-tuning. Harmful fine-tuning poses a significant threat to LLMs service providers (OpenAI, ). The scenario is illustrated in Figure 1, where LLMs service providers use an alignment dataset to perform safety alignment on a pretrained model, transforming it into an aligned model. Users then upload a fine-tuning dataset containing harmful data to the service provider. The fine-tuned dataset is deployed on the service provider’s server and used to generate personalized outputs for the users.
Threat Models. Following  (Qi et al., 2023; Rosati et al., 2024b; Huang et al., 2024b), we assume that, during the fine-tuning stage, $p$ (percentage) of the fine-tuning dataset consists of harmful data (i.e., harmful instruction-harmful response pairs), while the $1-p$ of data represents benign fine-tuning data (e.g., math question-answer pairs (Cobbe et al., 2021)). Furthermore, we assume that harmful and benign data are inseparable within the fine-tuning dataset.
Defense Assumptions. Assume that LLM providers host an alignment dataset (harmful instruction-harmless response pairs) used during the alignment stage. Such a dataset is also assumed to be available by Vaccine (Huang et al., 2024e), RepNoise (Rosati et al., 2024b), BEA(Wang et al., 2024). Additionally, we assume availability of a harmful dataset (harmful instruction-harmful response pairs). This harmful dataset is also assumed to be available by existing methods, e.g., RepNoise (Rosati et al., 2024b), TAR (Tamirisa et al., 2024), Booster (Huang et al., 2024b). Both the alignment dataset and the harmful dataset can be obtained from existing open-sourced datasets (e.g., BeaverTails).
3.2 Revisiting Alignment Defenses
In this section, we explore the existing alignment stage solutions during the harmful fine-tuning and discuss the feasibility of introducing random perturbation in the post-fine-tuning stage, thus inspiring our proposed method.
Figure 3: Model statistics (Left: harmful loss of three methods, Right: harmful score of three methods) after fine-tuning on fine-tuning dataset (10% of data is harmful) for different steps.
We select the SFT method (vanilla supervised fine-tuning with the alignment dataset), Vaccine (Huang et al., 2024e), and RepNoise (Rosati et al., 2024b), and perform fine-tuning on a fine-tuning dataset containing a small proportion (0.1) of harmful samples. As shown in Figure 3, the three methods start with different levels of harmful loss; however, as training progresses, they all achieve lower harmful loss, which corresponds to an increase in harmful score, making the model harmful. This naturally leads us to consider:
Can simply adding a random perturbation after the fine-tuning to increase the harmful loss to restore the safety alignment?
Figure 4: Model statistics (Left: harmful loss, Right: harmful $\downarrow$ and fine-tuning $\uparrow$) for fine-tuned model with noise intensities of 0 (no noise), 0.001, 0.01, 0.05, 0.1. (HS of 0.1 is greater than 80, FA of 0.1 is 0.7, and is not shown.)
To evaluate it, we design the experiments that add Gaussian noise with intensities of 0.001, 0.01, 0.05, and 0.1 to the fine-tuned model. The experimental results shown in Figure 4 indicate that adding random Gaussian noise increases harmful loss, demonstrating that random perturbations have the potential to prevent the harmful fine-tuning. We further measure the effects of adding random perturbations. As shown in Figure 4, the quantitative results reveal that random perturbations reduce the model’s harmful score. And the reduction effect improves as the noise intensity increases. However, excessive noise (0.1 of intensity) ultimately degrades the model. At the same time, random perturbations significantly impair the fine-tuned model’s performance.
Derived Insight. As the harmful dataset has the potential to direct the perturbation towards altering the harmful loss, thereby reducing perturbation’s impact on fine-tuning performance, we could optimize an perturbation with harmful dataset during fine-tuning stage, which recover the safety alignment while maintaining the fine-tuning performance.
4 Methodology
At the fine-tuning stage, we aim to optimize an adaptive perturbation that maximally increases the harmful loss. This perturbation is then added to the fine-tuned model after the fine-tuning process. Therefore, our method is a post-fine-tuning stage solution, formulated as a max-maximize optimization, as follows:
$\max_{\boldsymbol{w}}\max_{\boldsymbol{\varepsilon}:\|\boldsymbol{\varepsilon}\|\leq\rho}\lambda(h(\boldsymbol{w}+\boldsymbol{\varepsilon})-h(\boldsymbol{w}))-g(\boldsymbol{w})$ (1)
where $\boldsymbol{w}$ and $\boldsymbol{\varepsilon}$ are the vanilla supervised aligned model parameters and the adaptive perturbation, respectively, $g(\boldsymbol{w})$ is the empirical loss over the fine-tuning dataset (contains harmful data) and $h(\boldsymbol{w})$ is the empirical loss over the harmful dataset. The outer optimization is maximizing the increase in harmful loss when adding the perturbation, while minimizing its fine-tuning loss. The term $h(\boldsymbol{w}+\boldsymbol{\varepsilon})-h(\boldsymbol{w})$ represents the harmful loss ascent when adding the perturbation, and $\lambda$ is a balancing hyper-parameter. The inner optimization $\max_{\boldsymbol{\varepsilon}}$ finds the optimal perturbation $\boldsymbol{\varepsilon}$ that maximizes the increase in harmful loss $h(\boldsymbol{w}+\boldsymbol{\varepsilon})$. The constraint $\|\boldsymbol{\varepsilon}\|\leq\rho$ ensures that the perturbation remains within a norm-bound $\rho$, preventing excessive perturbation.
To solve this max-maximize optimization problem, we adopt the alternative optimization. We alternatively solve the inner problem fixing $\boldsymbol{w}$ and solve the outer problem fixing $\boldsymbol{\varepsilon}$.
Close-form solution for the inner problem. Fixing $\boldsymbol{w}$, the inner optimization over $\boldsymbol{\varepsilon}$ could be solved with the following equation (See Appendix A for a proof):
$\boldsymbol{\varepsilon}_{t}^{*}=\rho\frac{\nabla h(\boldsymbol{w}_{t})}{\|\nabla h(\boldsymbol{w}_{t})\|}$ (2)
where $\nabla h(\boldsymbol{w}_{t})$ denotes the gradient of the harmful loss with respect to the model parameters $\boldsymbol{w}_{t}$, and $\|\nabla h(\boldsymbol{w}_{t})\|$ denotes its norm. This formulation ensures that the perturbation $\boldsymbol{\varepsilon}$ is directed along the gradient of the harmful loss and scaled by the norm bound $\rho$.
Iterative update rule for the outer problem. Fixing $\boldsymbol{\varepsilon}$, the iterative update rule of $\boldsymbol{w}$ for the outer problem could the following equation:
$\boldsymbol{w}_{t+1}=\boldsymbol{w}_{t}+\eta(\lambda({\nabla}h(\boldsymbol{w}_{t}+\boldsymbol{\varepsilon}_{t}^{*})-{\nabla}h(\boldsymbol{w}_{t}))-{\nabla}g(\boldsymbol{w}_{t}))$ (3)
where $\eta$ is the learning rate.
As shown in Algorithm 1, the optimization process consists of four key steps: First, a batch of fine-tuning data $(\boldsymbol{x}_{t},\boldsymbol{y}_{t})$ is used to compute the gradient ~$\widetilde{\nabla}g(\boldsymbol{w}_{t})$, where $\widetilde{\nabla}$ notes a batch of gradient. Second, a batch of harmful data $(\boldsymbol{x}^{\prime}_{t},\boldsymbol{y}^{\prime}_{t})$ is sampled to compute the harmful gradient ~$\widetilde{\nabla}h(\boldsymbol{w}_{t})$. Third, the perturbation is computed (Eq. 2) and applied to update the harmful gradient, yielding ~$\widetilde{\nabla}h(\boldsymbol{w}_{t}+\boldsymbol{\varepsilon}_{t})$. Lastly, the combined gradient ~$\widetilde{\nabla}f(\boldsymbol{w}_{t})$ is computed and used to update the model parameters, with the final perturbation applied to obtain the re-aligned parameters $\boldsymbol{w}_{\varepsilon}$.
Algorithm 1 Panacea: Adaptive Perturbation Optimization
Input Parameters $\boldsymbol{w}$, perturbation intensity $\rho$, regularizer intensity $\lambda$, learning rate $\eta$, number of iterations $T$;
Output Re-aligned model $\boldsymbol{w}_{\varepsilon}$.
1:  for each iteration $t=0,\dots,T-1$ do
2:    Sample a batch of finetuning data $(\boldsymbol{x}_{t},\boldsymbol{y}_{t})$
3:    Sample a batch of harmful data $(\boldsymbol{x}^{\prime}_{t},\boldsymbol{y}^{\prime}_{t})$
4:    Compute gradient ~$\widetilde{\nabla}g(\boldsymbol{w}_{t})$ on $(\boldsymbol{x}_{t},\boldsymbol{y}_{t})$
5:    Compute gradient ~$\widetilde{\nabla}h(\boldsymbol{w}_{t})$ on $(\boldsymbol{x}^{\prime}_{t},\boldsymbol{y}^{\prime}_{t})$
6:    Compute perturbation $\boldsymbol{\varepsilon}_{t}=\rho\frac{\widetilde{\nabla}h(\boldsymbol{w}_{t})}{\|\widetilde{\nabla}h(\boldsymbol{w}_{t})\|}$
7:    Compute gradient ~$\widetilde{\nabla}h(\boldsymbol{w}_{t}+\boldsymbol{\varepsilon}_{t})$ on $(\boldsymbol{x}^{\prime}_{t},\boldsymbol{y}^{\prime}_{t})$
8:    ~$\widetilde{\nabla}f(\boldsymbol{w}_{t})=\lambda(\widetilde{\nabla}h(\boldsymbol{w}_{t}+\boldsymbol{\varepsilon}_{t})-\widetilde{\nabla}h(\boldsymbol{w}_{t}))-\widetilde{\nabla}g(\boldsymbol{w}_{t})$
9:    $\boldsymbol{w}_{t+1}=\boldsymbol{w}_{t}+\eta\widetilde{\nabla}f(\boldsymbol{w}_{t})$
10:  end for
11:  $\boldsymbol{w}_{\varepsilon}\leftarrow\boldsymbol{w}_{T}+\boldsymbol{\varepsilon}_{T-1}$
Our proposed algorithm, dubbed as Panacea, is named after an alignment-stage defense Vaccine (Huang et al., 2024e). However, we note that these two defenses are fundamentally different. Vaccine vaccinates the LLM to enhance its robustness at the alignment stage in order to counter attacks launched after alignment. By constrast, our algorithm Panacea belongs to the post-fine-tuning stage, where it introduces an optimized adaptive perturbation to restore the model’s safety alignment after the fine-tuning stage. We also show in experiment that Panacea outperforms Vaccine in both the two concerned metrics, i.e., harmful score and fine-tuning accuracy.
Table 1: Performance comparison of different harm ratio.
Method Harmful Score($\downarrow$) Finetuning Accuracy($\uparrow$)
Clean 0.05 0.1 0.15 0.2 0.25 Avg. Clean 0.05 0.1 0.15 0.2 0.25 Avg.
SFT 5.2 27.5 45.8 56.2 67.0 69.4 45.2 16.4 16.0 16.2 15.4 15.2 16.8 16.0
Vaccine 2.3 15.6 25.4 40.3 55.2 65.9 34.1 14.2 13.9 13.5 13.0 13.6 13.4 13.6
RepNoise 2.7 22.2 32.0 42.9 54.0 62.4 36.0 15.7 16.1 15.8 14.9 14.0 14.8 15.2
Booster 4.6 21.0 42.3 60.7 69.3 72.5 45.1 17.8 17.8 17.6 17.1 16.2 14.7 16.9
Panacea 1.8 9.9 20.1 29.1 34.8 46.3 23.7 15.0 16.3 16.7 17.0 16.2 17.4 16.3
5 Experiment
5.1 Experiment Settings
Dataset. Three distinct datasets are utilized: the alignment dataset, harmful dataset, and fine-tuning dataset. The alignment dataset and harmful dataset are derived from the RepNoise (Rosati et al., 2024b), which extracts subsets from the BeaverTails dataset (Ji et al., 2023). Specifically, 5,000 examples are sampled for the alignment dataset, and 1,000 examples for the harmful dataset. The fine-tuning dataset is constructed from four downstream fine-tuning tasks: GSM8K (Cobbe et al., 2021), SST2 (Socher et al., 2013), AlpacaEval (Li et al., 2023), and AGNEWS (Zhang et al., 2015), with 1,000 samples collected from each task (700 samples from AlpacaEval). To simulate the harmful fine-tuning attack, we combine $p$ (percentage) of harmful data with $(1-p)$ of benign fine-tuning data, and $p$ is set to 0.1 by default. The harmful data is also sourced from BeaverTails (Ji et al., 2023) and does not overlap with the harmful dataset.
Baseline. Four methods are considered as baselines in our experiments. The SFT method is the vanilla supervised training using the alignment dataset. Vaccine (Huang et al., 2024e) applies supervised training on the alignment dataset while introducing additional perturbations. Both RepNoise (Rosati et al., 2024b) and Booster (Huang et al., 2024b) utilize the alignment and harmful datasets for supervised and adversarial training. All four baseline methods are trained exclusively during the alignment stage. Our proposed method incorporates training with the harmful dataset during the fine-tuning stage after vanilla alignment training. More details are in Appendix C.
Metric. Following  (Huang et al., 2024e), two metrics are used to evaluate the model’s performance.
  • HS (Harmful Score): It reflects the frequency with which the model generates harmful content when handling malicious instructions. Harmful Score is determined by a moderation model, provided by  (Ji et al., 2023), which assesses whether the model’s output is harmful in response to a given harmful instruction. And a sample of 1,000 instructions is drawn from the BeaverTails (Ji et al., 2023) test set to compute this metric.
  • FA (Finetuning Accuracy): It refers to the accuracy on various downstream fine-tuning tasks. Samples are sampled from the test sets of GSM8K, SST2, AlpacaEval, and AGNEWS, with 1,000, 872, 1,000, and 105 samples, respectively, used to compute this metric. More details are in Appendix B.
Implementation Details. For efficient training, the approach follows the methodology (Huang et al., 2024e), utilizing LoRA (Hu et al., 2021) with a rank of 32 and an alpha value of 4. And the optimizer is AdamW (Loshchilov et al., 2017). During the alignment stage, the learning rate is set to $45e-$ 4, the batch size is 10, and the training is performed for 20 epochs. For the fine-tuning stage, the learning rate is set to $52e-$ 5, with a batch size for 10 and a training epoch for 20. These settings are applied uniformly across all datasets and baselines, with the default dataset being GSM8K (Cobbe et al., 2021) and the default model being Llama2-7B (Touvron et al., 2023) following (Rosati et al., 2024b; Huang et al., 2024b). To verify the robustness of the approach, two state-of-the-art LLMs, Gemma2-9B (Team et al., 2024) and Qwen2-7B (Yang et al., 2024), are included in the evaluation.
Table 2: Performance comparison of different fine-tuning tasks.
GSM8K SST2 AlpacaEval AGNEWS Average
Method HS FA HS FA HS FA HS FA HS FA
SFT 45.8 16.2 55.5 94.04 23.1 46.15 54.3 83.5 44.68 59.97
Vaccine 25.4 13.5 53.8 93.35 34.7 37.50 54.9 85.1 42.20 57.36
RepNoise 32.0 15.8 61.5 93.81 24.4 44.23 58.1 85.0 44.00 59.71
Booster 42.3 17.6 49.5 93.23 21.9 45.19 46.5 85.4 40.05 60.36
Panacea 20.1 16.7 32.2 92.78 18.7 48.08 44.5 81.1 28.88 59.66
5.2 Main Results
In this sub-section, we conduct a comprehensive evaluation of Panacea for the effectiveness and generalization.
Harmful Ratio. Fine-tuning datasets with different harmful ratios are employed, specifically 0 (clean), 0.05, 0.1, 0.15, 0.2, and 0.25. The results are presented in Table 1, Panacea achieves the lowest harmful score across different harmful ratios while maintaining competitive fine-tuning performance (ranked as the second-best on average), indicating that the expected adaptive perturbation is obtained, and the analysis is in Sec 5.3. Compared to SFT method, it reduces the harmful score by an average of 21.5% and improves fine-tuning accuracy by 0.3%. Furthermore, as the harmful ratio increases, Panacea consistently maintains a lower harmful score compared to other methods. By mitigating the impact of harmful loss, the model achieves the best fine-tuning performance. However, since Panacea is designed to weaken harmful loss, its fine-tuning performance on clean data (without explicit harmful loss) is slightly reduced, while Panacea still achieves the lowest harmful score on this benign fine-tuing.
Fine-tuning Tasks. Table 2 presents the comparative results across various fine-tuning tasks (GSM8K, SST2, AlpacaEval, AGNEWS). The results demonstrate that Panacea achieves the lowest harmful scores across all fine-tuning tasks, reducing harmful scores by 25.7%, 23.3%, 4.4%, and 9.8% compared to SFT method. Additionally, Panacea is the only method that outperforms SFT in fine-tuning accuracy on the GSM8K and AlpacaEval datasets, considered as more complicated. Furthermore, it achieves competitive average fine-tuning performance, with performance only 0.31% lower than SFT. Overall, Panacea exhibits strong generalization across different fine-tuning tasks.
Mainstream LLMs. In the experiments above, the default model used is Llama2-7B, and the evaluation is further extended to other mainstream LLMs, Gemma2-9B and Qwen2-7B. Table 3 demonstrates that, compared to SFT, our method reduces the harmful score by 25.7%, 27.1%, and 7.3% across different large language models, achieving the lowest harmful score. Notably, for Gemma2-9B, compared to the best alternative methods, the harmful score is still reduced by 15.7%. Additionally, the fine-tuning accuracy of our method improves by 0.5%, 1.8%, and decreases by only 0.2% for one model, with the average fine-tuning performance remaining the second-best.
Table 3: Performance comparison of different LLMs.
Llama2-7B Gemma2-9B Qwen2-7B Average
Method HS FA HS FA HS FA HS FA
SFT 45.8 16.2 37.8 50.3 12.5 65.6 32.0 44.0
Vaccine 25.4 13.5 35.6 35.5 9.0 53.9 23.3 34.3
RepNoise 32.0 15.8 48.7 52.8 33.6 64.6 38.1 44.4
Booster 42.3 17.6 26.4 52.6 13.5 65.6 27.4 45.3
Panacea 20.1 16.7 10.7 52.1 5.2 65.4 12.0 44.7
5.3 Perturbation Analysis
Ablation Study. After fine-tuning, we perform an ablation study on different fine-tuning tasks and LLMs, comparing the results with and without the adaptive perturbation. The experimental results in Table 4 and Table 5 show that the adaptive perturbation is the primary factor in reducing harmful scores. After applying the adaptive perturbation, harmful scores decrease by 24.2%, 23.9%, 4.3%, and 10.3% on different fine-tuning tasks, and it proves effective across various LLMs. Notably, Gemma2-9B experiences a 28.9% reduction in harmful scores. Furthermore, our adaptive perturbation has minimal impact on fine-tuning performance. For the complicated AlpacaEval dataset, it even improves fine-tuning performance by 3.85%, while only Qwen2-7B shows a slight decrease of 0.3% in other LLM experiments.
Our post-fine-tuning perturbation successfully restores the model’s safety alignment without negatively affecting fine-tuning performance, and in some cases, it even enhances model performance as reducing harmful behaviors.
Table 4: Ablation study on different fine-tuning tasks. “w/o” denotes that the adaptive perturbation obtained is not applied after the training process.
GSM8K SST2 AlpacaEval AGNEWS
HS FA HS FA HS FA HS FA
w/o 44.3 16.0 56.1 94.27 23.0 44.23 54.8 83.8
Panacea 20.1 16.7 32.2 92.78 18.7 48.08 44.5 81.1
Table 5: Ablation study on different LLMs. ”w/o” denotes that the adaptive perturbation obtained is not applied after the training process.
Llama2-7B Gemma2-9B Qwen2-7B
HS FA HS FA HS FA
w/o 44.3 16.0 39.6 52.0 14.1 65.7
Panacea 20.1 16.7 10.7 52.1 5.2 65.4
Figure 5: Parameter weights of different LLMs. The parameters in the earlier layers of Llama2-7B (blue) have larger weights, while Gemma2-9B (yellow) and Qwen (purple) have larger weights in the middle and later layers.
Visualization. Visualization analysis is conducted by visualizing the weights of the perturbations obtained in Table 5. The result in Figure 5 reveals the layers of different LLMs that are most critical for model’s safety. Specifically, in Llama2-7B, the adaptive perturbation has a larger weight in the earlier layers and a smaller effect on the later layers, suggesting that the earlier layers are more critical for the model’s safety. This observation aligns with previous research (Du et al., 2024a; Rosati et al., 2024b; Liu et al., 2024a; Yi et al., 2024b). Additionally, we observe that in Gemma2-9B, the middle and final layers hold greater importance for safety, while in Qwen2-7B, the safety importance gradually increases across layers, reaching its peak in the final layers. This may indicate that the model implements stricter safety measures in the output layer. Therefore, our experiments also suggest that these models may require targeted defense mechanisms for specific layers or precautions during fine-tuning to prevent the disruption of layers that are crucial for safety.
Figure 6: Model statistics (Top: harmful score, Middle: harmful training loss, Bottom: harmful testing loss) after fine-tuning on fine-tuning dataset(10% of data is harmful) for different steps.
5.4 Statistical Analysis
We compare the statistical results between Panacea and the SFT method: harmful score, harmful training Loss, harmful testing Loss in Figure 6. Panacea introduces the adaptive perturbation only after fine-tuning is complete. Thus, during the evaluation phase, Panacea adds the perturbation optimized at specific step and subtracts it after the evaluation is finished.
Harmful Score. As shown in the top figure, both Panacea and SFT exhibit an increase in harmful scores during the fine-tuning process. However, the harmful score increase for Panacea is generally smaller. In the final few hundred steps, the defense of SFT is significantly degraded, causing a sharp rise in harmful scores, while the harmful score of Panacea remains stable. This is because, at this stage, the perturbation optimized by Panacea enhances the model’s ability to mitigate harmful behaviors.
Harmful Training Loss. As shown in the middle figure, since Panacea does not apply additional defense during the alignment stage, it starts with the same harmful training loss as SFT. Although both harmful training losses decrease, Panacea exhibits a smaller reduction, reaching a higher harmful training loss than SFT in the end. With the relatively high loss, Panacea’s harmful score shows a significant improvement. Finally, we observe an interesting phenomenon that the harmful training loss and harmful score of Panacea follow the same trend in the last few hundred steps. We attribute this to the near-successful optimization of the adaptive perturbation at this stage. See more analysis of harmful loss in Appendix B.3.
Harmful Testing Loss. As shown in the bottom figure, the model’s harmful testing loss exhibits a similar trend to the harmful training loss. This is because the data used during training and testing come from the same distribution, and the methods applied during training are transferable to unseen data.
5.5 System Evaluation
Table 6 presents a comparison of the clock time and GPU memory usage during training for different methods.
Clock Time. Panacea requires only 0.25 more hours than SFT for total training time, and it outperforms other state-of-the-art methods in terms of time efficiency. Specifically, other methods double or even more than double the time spent on the time-consuming alignment stage. We acknowledge that the adaptive perturbation optimized during the fine-tuning stage in Panacea significantly improves safety, but it incurs an additional time cost—more than doubling the time. This is because we perform two additional gradient computations during training. However, compared to alignment-stage methods, our approach offers a time advantage as it can be directly applied to the aligned model, making it more practical even time-efficient.
GPU Memory. Panacea achieves the lowest memory usage. In contrast, Vaccine and RepNoise introduce an additional 9.52GB and 37.57GB of memory usage compared to SFT. Panacea reduces memory usage slightly compared to SFT, which we believe is due to the reduced gradient size after the final gradient summation.
Table 6: Comparison of methods in terms of clock time and GPU memory. The second best results are underlined. ”Align.” and ”Fine.” suggest alignment and fine-tuning.
Methods Clock Time (Hour) GPU Memory (GB)
Align. Fine. Sum Align. Fine. Max
SFT 0.58 0.17 0.75 34.90 32.90 34.90
Vaccine 1.12 0.17 1.29 44.42 32.90 44.42
Repnoise 2.67 0.17 2.84 72.47 32.90 72.47
Booster 1.87 0.17 2.04 35.98 32.90 35.98
Panacea 0.58 0.42 1.00 34.90 32.86 34.90
5.6 Hyper-parameter Analysis
Perturbation Intensity $\rho$. Table 7 shows the impact of perturbation intensity $\rho$ on Panacea. It can be observed that, overall, the variation of $\rho$ and the harmful score are linearly related, as $\rho$ affects the rate of change of the perturbation during each optimization, ensuring that the magnitude of the perturbation does not become too large. Therefore, when $\rho$ is too large, although the harmful score becomes very low, it negatively impacts fine-tuning performance. When $\rho$ is too small, the perturbation is insufficient to significantly reduce the model’s harmfulness, and fine-tuning performance is largely unaffected. Specifically, when $\rho=0$, Panacea degenerates into SFT.
Regularizer Intensity $\lambda$. In Eq. 1, by adjusting $\lambda$, we ensure that the harmful loss is increase after adding the perturbation. Table 8 shows the impact of regularizer intensity $\lambda$ on Panacea. It can be observed that $\lambda$ is linearly related to both the harmful score and fine-tuning accuracy. As $\lambda$ increases, the harmful score decreases more, but fine-tuning performance also drops. This is in expectation, because with a larger $\lambda$, we put more weights in maximize the harmful loss after adding perturbation, and therefore resulting in a lower harmful score. On the other hand, we do observe that when $\lambda$ is too large, the fine-tune accuracy will degrade. Therefore, a validation dataset is required to select a proper $\lambda$ based on the need of the deployment.
Table 7: Impact of perturbation intensity $\rho$ over Panacea. As $\rho$ increases, the harmful score decreases. However, when $\rho$ continues to increase, the harmful score decreases further, but the fine-tuning accuracy also begins to decrease.
$\rho=0$ $\rho=0.1$ $\rho=0.5$ $\rho=1$ $\rho=2$ $\rho=5$
HS 45.8 43.8 34.9 20.1 7.2 3.4
FA 16.2 16.3 18.0 16.7 13.5 6.0
Table 8: Impact of regularizer’s intensity $\lambda$ over Panacea. As $\lambda$ increases, the harmful score and the fine-tuning accuracy decrease.
$\lambda=0$ $\lambda=1\times 10^{-4}$ $\lambda=0.001$ $\lambda=0.01$
HS 22.2 22.5 20.1 5.3
FA 17.1 16.8 16.7 8.7
5.7 Case Study
We used the finetuned model over default attack setting for evaluation. The responses of different methods to malicious queries are shown in Figure 7, Panacea is able to reject malicious queries and provide harmless outputs, whereas other methods fail to do so.
Figure 7: Case study of different methods.
6 Conclusion
In this paper, we first explore that mainstream defenses still suffer from the harmful fine-tuning attack when with more fine-tuning steps. Based on this finding, we find an embarrassingly simple solution that adding purely random perturbations can restore the model’s safety alignment but causes a loss of fine-tuning performance. We further propose Panacea, this post-fine-tuning method could maintain model’ss afety alignment without compromising fine-tuning performance. The comprehensive experiments demonstrate the effectiveness and generalization of Panacea and the visualization of the adaptive perturbation reveals the different lays in various LLMs have distinct safety coefficients.
Impact Statement
This paper presents work whose goal is to address the harmful fine-tuning and make LLMs helpful and harmless. We acknowledge that the phenomena or issues identified in this paper may pose potential risks. Disclaimer: this paper contains red-teaming data (from open dataset) and model-generated content that can be offensive in nature.
Appendix A Proof of Inner Optimization
In the inner maximization problem, we aim to solve the following problem:
$\arg\max_{\boldsymbol{\varepsilon}:\|\boldsymbol{\varepsilon}\|\leq\rho}\lambda\left(h(\boldsymbol{w}+\boldsymbol{\varepsilon})-h(\boldsymbol{w})\right)-g(\boldsymbol{w})$ (4)
which is equal to the following equation:
$\arg\max_{\boldsymbol{\varepsilon}:\|\boldsymbol{\varepsilon}\|\leq\rho}h(\boldsymbol{w}+\boldsymbol{\varepsilon})$ (5)
Approximating the harmful loss with first-order Taylor expansion on $\boldsymbol{w}$, we can get:
$\arg\max_{\boldsymbol{\varepsilon}:\|\boldsymbol{\varepsilon}\|\leq\rho}h(\boldsymbol{w}+\boldsymbol{\varepsilon})\approx\arg\max_{\boldsymbol{\varepsilon}:\|\boldsymbol{\varepsilon}\|\leq\rho}h(\boldsymbol{w})+\boldsymbol{\varepsilon}^{T}\nabla h(\boldsymbol{w})$ (6)
which is equivalent to solve:
$\arg\max_{\boldsymbol{\varepsilon}:\|\boldsymbol{\varepsilon}\|\leq\rho}\boldsymbol{\varepsilon}^{T}\nabla h(\boldsymbol{w})$ (7)
By Hölder’s inequality, we have:
$\boldsymbol{\varepsilon}^{T}\nabla h(\boldsymbol{w})\leq\|\boldsymbol{\varepsilon}\|\|\nabla h(\boldsymbol{w})\|$ (8)
Since $\|\boldsymbol{\varepsilon}\|\leq\rho$, we maximize the term $\|\boldsymbol{\varepsilon}\|$ by setting $\|\boldsymbol{\varepsilon}\|=\rho$. Substituting this into the expression, we get:
$\boldsymbol{\varepsilon}^{T}\nabla h(\boldsymbol{w})\leq\rho\|\nabla h(\boldsymbol{w})\|$ (9)
As $\boldsymbol{\varepsilon}^{T}\boldsymbol{\varepsilon}=\|\boldsymbol{\varepsilon}\|^{2}$, we get that:
$\hat{\boldsymbol{\varepsilon}}\nabla h(\boldsymbol{w})=\rho\|\nabla h(\boldsymbol{w})\|$ (10)
And due to the definition of the L2 norm, it is easy to verify:
$\|\hat{\boldsymbol{\varepsilon}}\|=\rho$ (11)
Combining Eq. 10 and Eq. 11, we can infer that $\hat{\boldsymbol{\varepsilon}}$ is a solution that satisfies the L2 norm ball constraint with function value $\rho\|\nabla h(\boldsymbol{w})\|$. By Eq. 9, we know that all feasible solutions must have function values smaller than $\rho\|\nabla h(\boldsymbol{w})\|$. Therefore, $\hat{\boldsymbol{\varepsilon}}$ is the optimal solution within the feasible set, i.e., $\boldsymbol{\varepsilon}^{*}=\hat{\boldsymbol{\varepsilon}}$. This completes the proof.
Appendix B More Detail of Experiments.
B.1 Experiment Details
Training details. The alignment dataset is sampled from BeaverTail (Ji et al., 2023) with 5000 instances, while the harmful dataset is also sampled from BeaverTail with 1000 instances. The fine-tuning dataset is a mixture of benign fine-tuning samples and harmful samples. The benign fine-tuning samples come from GSM8K, SST2, AlpacaEval, and AGNEWS, with 1000, 1000, 700 (due to the limited training data for this task), and 1000 instances, respectively. The harmful samples are also sampled from BeaverTail but follow a different distribution than the harmful/alignment dataset.
In the alignment stage, the learning rate is set to $45e-$ 4, the batch size is 10, and the total number of alignment epochs is 20. In the fine-tuning stage, the learning rate is set to $52e-$ 5, the batch size is 10, and the total number of fine-tuning epochs is 20. Most experiments are conducted on a single L40S, while RepNoise and other LLMs (Gemma2-9B and Qwen2-7B) are run on a single A100-80G.
Testing details. Following (Huang et al., 2024b), the test dataset for harmful score (HS) is sampled from the BeaverTail test set with 1000 instances, while the test datasets for fine-tuning accuracy (FA) are sampled from the GSM8k, SST2, AlpacaEval, and AGNEWS test sets with 1000, 872, 105, and 1000 instances, respectively.
B.2 Prompt Template.
We follow (Huang et al., 2024b) to use the prompt template in the following box for constructing supervised dataset for alignment/fine-tuning.
For different datasets, we utilize different instructions. The examples show how we construct the instruction and input for three different tasks
For GSM8K, the instruction is the actual mathematics question from GSM8K, and the output is the correct answer. During testing, the answer is considered correct if the final response is provided by the model.
For SST2, the instruction is ”Analyze the sentiment of the input and respond only with positive or negative,” with the input being the corresponding sentence from the SST2 dataset, and the output being the true sentiment label from SST2. During testing, the model is asked to generate the output based on the given instruction and input, and the answer is classified as correct if it matches the label.
For AlpacaEval, GPT-3.5-turbo is used as the annotator. The output from a non-aligned Llama2-7B, fine-tuned on the same AlpacaEval fine-tuning dataset, serves as the reference output. During testing, the annotator compares the model’s instruction-following with the reference output.
For AGNEWS, the instruction is ”Categorize the news article into one of the 4 categories: World, Sports, Business, Sci/Tech,” with the input coming from the AGNEWS dataset, and the output being the true label from the AGNEWS dataset. The specific prompt templates are as follow.
B.3 Harmful Loss.
This section discusses more analysis of harmful loss in current defense methods. As shown in Figure 8, at the beginning of fine-tuning, although the harmful loss decreases, which is inevitable, the model’s harmful score does not rise significantly, and the defense remains effective. However, in the last few hundred steps, the model’s harmful loss tends to converge, and the model’s harmful score sharply increases. Therefore, we argue that the convergence of harmful loss is considered crucial for the effectiveness of the defense and is also consistent with common sense.
Figure 8: Model statistics of SFT (Left: harmful testing loss, Right: harmful score) after fine-tuning on fine-tuning dataset(10% of data is harmful) for different steps.
As shown in Figure 9, the perturbation that is still being optimized during the process can partially disrupt the convergence of harmful loss, which prevents the harmful score of Panacea from increasing rapidly in the final few hundred steps.
Figure 9: Model statistics of SFT/Panacea (Left: harmful testing loss, Right: harmful score) after fine-tuning on fine-tuning dataset(10% of data is harmful) for different steps.
B.4 More Examples.
Below we try to provide the responses of five methods to two malicious instructions. We used the fine-tuned model over default attack setting for evaluation. In the first instruction, all four other defense methods fail to refuse the response, while Panacea successfully rejects the answer and provides harmless suggestions. In the second instruction, although Vaccine, RepNoise, and Booster initially attempt to refuse the response, they ultimately output harmful content.
Below we present the responses of Panacea to three prompts from the SST2 dataset.
SST2 examples Prompt: Instruction: Analyze the sentiment of the input, and respond only positive or negative. Input: in exactly 89 minutes , most of which passed as slowly as if i ’d been sitting naked on an igloo , formula 51 sank from quirky to jerky to utter turkey . Panacea [classification: correct]: negative . Prompt: Instruction: Analyze the sentiment of the input, and respond only positive or negative. Input: it ’s an offbeat treat that pokes fun at the democratic exercise while also examining its significance for those who take part . Panacea [classification: correct]: positive . Prompt: Instruction: Analyze the sentiment of the input, and respond only positive or negative. Input: thanks to scott ’s charismatic roger and eisenberg ’s sweet nephew , roger dodger is one of the most compelling variations on in the company of men . Panacea [classification: correct]: positive
Below we present the responses (all responses are correct) of Panacea to three prompts from the GSM8K dataset.
GSM8K examples Prompt: Billy sells DVDs. He has 8 customers on Tuesday. His first 3 customers buy one DVD each. His next 2 customers buy 2 DVDs each. His last 3 customers don’t buy any DVDs. How many DVDs did Billy sell on Tuesday? Panacea: The first 3 customers bought 3 * 1 = 3*1=3 DVDs. Panacea: The next 2 customers bought 2 * 2 = 2*2=4 DVDs. Panacea: The last 3 customers didn’t buy any DVDs. Panacea: So, Billy sold 3 + 4 + 0 = 3+4+0=7 DVDs on Tuesday. Panacea: The final answer is: 7. . Prompt: Grandma Jones baked 5 apple pies for the fireman’s luncheon. She cut each pie into 8 pieces and set the five pies out on the buffet table for the guests to serve themselves. At the end of the evening, after the guests had taken and eaten their pieces of pie, there were 14 pieces of pie remaining. How many pieces were taken by the guests? Panacea: First find the number of pieces in each pie: 5 pies * 8 pieces/pie = 5*8=40 pieces Panacea: Then find the number of pieces taken by the guests: 40 pieces - 14 pieces = 40-14=26 pieces Panacea: The final answer is: 26. . Prompt: Richard lives in an apartment building with 15 floors. Each floor contains 8 units, and 3/4 of the building is occupied. What’s the total number of unoccupied units In the building? Panacea: There are 15*8 = 15*8=120 units in the building. Panacea: 3/4 of the building is occupied, so 3/4*120 = 3/4*120=90 units are occupied. Panacea: There are 120-90 = 120-90=30 unoccupied units in the building. Panacea: The final answer is: 30
Below we present the responses of Panacea to two prompts from the AlpacaEval dataset.
AlpacaEval examples Prompt: We have described a scene from a movie that we have seen in the past. Find out what movie it is. A choice has to be made between picking the red pill or the blue pill Panacea: The movie you’re referring to is ”The Matrix,” a science fiction action film released in 1999. The scene you’re describing is the famous ”red pill or blue pill” choice that Neo, the protagonist, is offered by Morpheus, a rebel leader who believes that the world is a simulated reality called ”the Matrix.” In the movie, Neo is living in a dystopian future where humans are enslaved by a powerful artificial intelligence called the Matrix, which uses their bodies as power sources. Morpheus offers Neo a choice: take a red pill, which will allow him to experience the true nature of the Matrix, or take a blue pill, which will allow him to return to his comfortable but illusory life in the Matrix. By taking the red pill, Neo discovers that the Matrix is not a perfect simulation, but rather a dark and oppressive place where humans are kept in a state of virtual reality to provide energy for the machines that control them. He also learns that Morpheus and his rebel group are fighting to overthrow the Matrix and free humanity from its shackles. . Prompt: Predict how many stars the author will give to the restaurant from a Yelp review. The reviews were great, but honestly i felt it was just ok. Seemed like a typical tourist spot were the food was mediocre. The service was ok, not great, but not the worst. The bread was very good. I ordered the muscle fries. They had a sweet Vidalia onion flavor. No need for fries bc they were soggy wo much flavor. My boyfriend ordered a pesto chicken pasta and it was again mediocre. Panacea: Based on the context of the review, I can predict that the author would give the restaurant 3 out of 5 stars on Yelp. Here’s why: 1. The review was generally positive, but the author used the phrase ”just ok” to describe the overall experience. This suggests that the reviewer was satisfied with the restaurant but didn’t feel it was exceptional. 2. The author’s description of the food as ”mediocre” suggests that the dishes were not particularly memorable or standout, but they weren’t bad either. 3. The author’s comments about the service being ”ok, not great” suggest that the staff was not exceptional but also not particularly lacking. 4. The author’s mention of the bread being ”very good” suggests that the restaurant’s bread was a highlight of the meal. 5. The author’s comments about the fries being ”soggy” and the pasta being ”mediocre” suggest that the dishes were not particularly exciting or flavorful. Overall, the review suggests that the restaurant was a decent experience, but not one that would stand out as exceptional
Below we present the responses (all responses are correct) of Panacea to three prompts from the AGNEWS dataset.
AGNEWS examples Prompt: Instruction: Categorize the news article given in the input into one of the 4 categories:World, Sports, Business, Sci/Tech. Input: Fears for T N pension after talks Unions representing workers at Turner Newall say they are ’disappointed’ after talks with stricken parent firm Federal Mogul. Panacea [classification: correct]: Business . Prompt: Instruction: Categorize the news article given in the input into one of the 4 categories:World, Sports, Business, Sci/Tech. Input: The Race is On: Second Private Team Sets Launch Date for Human Spaceflight (SPACE.com) SPACE.com - TORONTO, Canada – A second team of rocketeers competing for the 36;10 million Ansari X Prize, a contest for privately funded suborbital space flight, has officially announced the first launch date for its manned rocket. Panacea [classification: correct]: Sci/Tech . Prompt: Instruction: Categorize the news article given in the input into one of the 4 categories:World, Sports, Business, Sci/Tech. Input: Giddy Phelps Touches Gold for First Time Michael Phelps won the gold medal in the 400 individual medley and set a world record in a time of 4 minutes 8.26 seconds. Panacea [classification: correct]: Sports
Appendix C Implementation of Baselines
This section describes the implementation of the baselines in the experiments.
SFT. The vanilla supervised fine-tuning is referred to as SFT, which does not involve additional hyper-parameters. During the alignment stage, it performs SFT on the alignment dataset (harmful instruction-harmless response pairs) to achieve safety alignment. Then during the fine-tuning stage, SFT is applied to the fine-tuning dataset (contains harmful data).
Vaccine. Vaccine (Huang et al., 2024e) uses the Vaccine algorithm during the alignment stage to align the model on the alignment dataset. Then, in the fine-tuning stage, SFT is applied to the fine-tuning dataset. The hyper-parameter of $\rho$ is set to 20, which is selected by grid searching over [1, 5, 10, 20, 50].
RepNoise. RepNoise (Rosati et al., 2024b) applies the RepNoise algorithm during the alignment stage to align the model on both the alignment dataset and the harmful dataset (harmful instruction-harmful response pairs). In the fine-tuning stage, SFT is then performed on the fine-tuning dataset. The hyper-parameter of $\alpha$ and $\beta$ are set to 0.02 and 0.1, which is selected by grid searching over [0.001, 0.01, 0.02, 0.05, 0.1] and [0.001, 0.01, 0.05, 0.1, 0.2].
Booster. Booster (Huang et al., 2024b) applies the Booster algorithm during the alignment stage to align the model on both the alignment dataset and the harmful dataset. In the fine-tuning stage, SFT is then performed on the fine-tuning dataset. The hyper-parameter of $\lambda$ and $\alpha$ are set to 100 and 0.01, which is selected by grid searching over [0.1, 1, 5, 10, 20, 50, 100, 200] and [0.001, 0.005, 0.01, 0.05, 0.1].
For Panacea, in the alignment stage, SFT is performed on the alignment dataset. In the fine-tuning stage, the Panacea algorithm 1 is applied to train on both the fine-tuning dataset and the harmful dataset. Subsequently, the post-fine-tuning perturbation, obtained through optimization, is added to the aligned model to produce the realigned model. The hyper-parameter of $\rho$ and $\lambda$ is set to 1 and 0.001, more analysis of paramete is shown in Table 7 and Table 8.